# Neuro-Symbolic Video Search

### v1

@@ 5.2 Ground Truth Temporal Logic Specifications
    Always/Eventually event A. These temporal logic queries can be expressed as either $\Phi=\Box\hskip 2.84526ptp_{k}$ or $\diamondsuit\hskip 2.84526ptp_{k}$, identifying occurrences of a single event within a video. For instance, a natural language equivalent of this query is “Identify all frames depicting a car acciden” or “Locate all frames showcasing an animal’s presence”.
@@
    Event A and B until C. This specification encapsulates a video where two simultaneously occurring events are monitored up to the point a third event occurs; for example “Recognize scenes that show a child playing and dog barking until an adult appears”. It is expressed by $\Phi=(p_{1}\wedge p_{2})\hskip 2.84526pt\mathsf{U}\hskip 2.84526ptp_{3}$.
The synthetic TLV dataset is constructed over all the above-mentioned TL specifications. In contrast, the AV datasets are more specifically focused on the TL specifications of “Event A until B” and “Event A and B until C”. This is due to the fact that generating some TL specifications for real datasets would require altering the original video sequences, a scenario we aim to avoid.
@@
Table 2: Comprehensive Performance Comparison Across Models and TLV datasets. This illustrates the impact of different neural perception models on the performance of NSVS-TL across various datasets. The meanings of each symbol are described in Sec. 5.2.
6 Experiment
  • Given the optimal perception model, how does the efficacy of NSVS-TL, which uses temporal logic to reason over detected propositions per frame, compare to that of a Large Language Model (LLM) with respect to different temporal logic specifications?
  • How does the duration of a video affect the performance of our method, particularly when the event is rare (in a long-duration video) or is temporally extended?
Neural Perception Models
We utilize various neural perception models to evaluate the performance of NSVS-TL. To assess the impact of perception models on scene identification, we leverage SOTA computer vision models such as YOLO V8 [29], Grounding Dino [1], Masked R-CNN [14], and Contrastive Language-Image Pre-training (CLIP) [28].
Benchmarks for Scene Identification
Analysis of our method, benchmarked against video-language foundation models such as Video-Llama and ViCLIP, reveals their inherent limitations in handling long-term and complex temporal queries (see Fig. 2). While these models perform well with shorter sequences, they lack robust temporal reasoning capabilities and struggle with extended video lengths. This limitation makes them less effective for long-term event detection tasks. Moreover, no current models exist that are capable of retrieving precise scenes or frames based on textual description or temporal logic specification, despite their ability to understand the semantics of video content. Hence, we design benchmarks for a balanced comparison that uses LLMs for reasoning over per-frame annotations instead of TL. Initially, neural perception models provide per-frame annotations to LLMs such as GPT. These LLMs are then prompted to identify scenes of interest according to specified propositions and TL specifications. We use these benchmarks to evaluate the impact of video length on scene identification performance. The prompts and methodologies applied in these comparisons are elaborated in the Appendix.
@@ 7 Results
Our work introduces a novel methodological approach that combines a neural perceptual model with symbolic methods. This integration provides a new perspective in understanding long-horizon videos. In fact, we propose a framework that allows for the integration of any neural perceptual models into our work, enhancing the capability to understand videos which enables us to localize frames of interest with respect to queries. We show a comparison between different neural perceptual models in Tab. 2 and Fig. 5(a). We observe that NSVS-TL with various neural perception models performs differently depending on the complexity of the TL specification and datasets. The variation underscores the importance of these models’ inherent visual detection capabilities for scene identification.
@@
We evaluate TL-based reasoning against LLM-based reasoning over per-frame semantic annotations derived from neural perception models. The evaluation is conducted on four datasets: COCO-TLV, ImageNet-TLV, NuScenes-TLV, and Waymo-TLV. The results are collectively presented in Fig. 5. For single event scenarios such as “Always/Eventually Event A Occurs”, both our method and LLM-based reasoning perform reasonably well since these events do not require complex reasoning. In multi-event scenarios (“Event A and Event B”, “Event A until Event B,” and “Event A and Event B until Event C”), our TL-based reasoning outperforms all LLM-based baselines.
@@
We evaluate multi-event sequences with temporally extended gaps, such as those in the specification of events A until B. For these scenarios, we design a scene identification task that extends the video length: event A starts at the beginning and continues until event B occurs at the end. This tests the reasoning capacity of the LLM benchmarks as the temporal distance between events increases. The scene identification performance of GPT3.5 and GPT3.5 Turbo Instruct degrade within videos that are $500\;s$ long as illustrated in Fig. 5(b). GPT4, on the other hand, suffers from a sharp decline in performance in videos whose lengths range from $1000\;s$ and more. In contrast, NSVS-TL maintains consistent performance throughout. This consistency is observed even in videos spanning up to $40$ minutes or $2400\;s$, indicating NSVS-TL’s potential reliability in handling even longer videos.
@@ 8 Conclusion
NSVS-TL, a neuro-symbolic method, enhances video understanding and scene identification in contexts demanding extended temporal reasoning. NSVS-TL, advances scene identification across various video durations by integrating semantic understanding with temporal reasoning. Unlike other methods that focus on short-duration video reasoning, we propose a novel approach for comprehending long-horizon, temporally extended videos.
@@
Supplementary Material
@@ 9.1 Examples of Temporal Logic
Note that compared to LTL, CTL can quantify multiple branches while LTL is capable of one path. However, LTL can encode formulas like ”Eventually-Always something happens,” while CTL is not capable of.
@@
LTL over infinite traces was introduced in the field of computer science as a specification language designed for concurrent programs [27]. It facilitates the characterization of temporal reasoning within the domains of propositions and first-order logic. In our approach, frames of interest are considered as a finite set of frames extracted from a video. We can employ Linear temporal logic over finite traces ($\text{LTL}_{f}$) [5, 33, 6, 3] to handle finite sequences encompassing finite-length trajectories, frames from video, system procedures. Given a set of atomic propositions $P$, the syntax of $\text{LTL}_{f}$ formulas is defined as follows:
@@ 9.2 LLM Benchmark Prompts
You are a Language Assistant to enable searching a video frame for objects satisfying a condition. TASK DESCRIPTION: I will be giving you the list of objects detected in the video and a condition from the search. The list of objects will be specified in this format and include those frames where an object is detected: - Frame 1: [object1, object2, object3] - Frame 2: [object1, object2] A condition is of the form: - object1 Until object2, - object1 Until (object2 and object3), - Eventually object 1, - Always object 2, etc. You can use the list of objects detected in the video to help you. You don’t have access to the video. Your job is to specify the frame numbers where the rule is satisfied. There can be many such sequences. If you cannot find any sequence, you must return output [None]. To clarify, you will be looking for frames where the rule is satisfied. The output is a list of frame numbers, such as [1,20, 25, 50], specifying the frames where the rule holds. An example of the list of detections per frame is as follows: - 01: [person, car] - 02: [person, car] - 10: [person, car] - 30: [truck] - 40: [car] - 44: [person] - 50: [car, truck] Case 1: ‘person’ until ‘truck’ gives all frames of ‘person’ and ‘truck’ if ‘truck’ shows up. Case 2: Eventually ‘person’ gives all frames of ‘person’ eventually showing up. Case 3: ‘car’ and ‘truck’ should give all frames of ‘car’, and ‘truck’ showing up in the same frame. Case 4: Always ‘car’ should give all frames of ‘car’ always showing up. Case 5: (‘car’ and ‘truck’) until ‘person’ gives all frames of ‘car’ and ‘truck’ showing up in the same frame if ’person’ shows up; all frames of ’person’ included.

### v3

@@ 5.2 Ground Truth Temporal Logic Specifications
    Always/Eventually event A. These temporal logic queries can be expressed as either $\Phi=\Box\hskip 2.84526ptp_{k}$ or $\diamondsuit\hskip 2.84526ptp_{k}$, identifying occurrences of a single event within a video. For instance, a natural language equivalent of this query is “Identify all frames depicting a car accident” or “Locate all frames showcasing an animal’s presence”.
@@
    Event A and B until C. This specification encapsulates a video where two simultaneously occurring events are monitored up to the point a third event occurs; for example, “Recognize scenes that show a child playing and dog barking until an adult appears”. It is expressed by $\Phi=(p_{1}\wedge p_{2})\hskip 2.84526pt\mathsf{U}\hskip 2.84526ptp_{3}$.
The synthetic TLV dataset is constructed based on all the TL specifications mentioned above. In contrast, the AV datasets are more specifically focused on the TL specifications of “Event A until B” and “Event A and B until C”. This is due to the fact that generating some TL specifications for real datasets would require altering the original video sequences, a scenario we aim to avoid.
6 Experiment
  • Given the optimal perception model, how does the efficacy of NSVS-TL, which employs TL to reason over detected propositions per frame, compare to that of a Large Language Model (LLM) across various TL specifications?
  • How does the duration of a video affect the performance of our method when the event is rare (in a long-duration video) or is temporally extended?
6.1 Neural Perception Models
We utilize various neural perception models to assess the impact of different perception models on scene identification. We leverage SOTA computer vision models such as YOLO V8 [25], Grounding Dino [39], Masked R-CNN [10], and Contrastive Language-Image Pre-training (CLIP) [24]. We use an image detection model instead of a video detection model for several reasons. First, we define a video as a sequence of frames, so detecting objects in each frame effectively means detecting objects throughout the entire video. Second, image detection can process per-frame input faster than video object detectors, making it essential for real-time video understanding, which is one of our key contributions.
6.2 Benchmarks for Scene Identification
Analysis of our method, benchmarked against video-language foundation models such as Video-Llama and ViCLIP, reveals their inherent limitations in handling long-term and complex temporal queries (see Fig. 2). While these models perform well with shorter sequences, they lack robust temporal reasoning capabilities and struggle with extended video lengths. This limitation makes them less effective for long-term event detection tasks. Moreover, no current models exist that are capable of retrieving precise scenes or frames based on textual description or TL specification, despite their ability to understand the semantics of video content. Hence, we design benchmarks for a balanced comparison that uses LLMs for reasoning over per-frame annotations instead of TL. Initially, neural perception models provide per-frame annotations to LLMs such as GPT. These LLMs are then prompted to identify scenes of interest according to specified propositions and TL specifications. We use these benchmarks to evaluate the impact of video length on scene identification performance. The prompts and methodologies applied in these comparisons are elaborated in the Appendix.
@@ 6.2 Benchmarks for Scene Identification
Table 2: Comprehensive Performance Comparison Across Models and TLV Datasets. This illustrates the impact of different neural perception models across various datasets. The meaning of each symbol is described in Sec. 5.2.
@@ 7 Results
Our work introduces a novel methodological approach that combines a neural perceptual model with symbolic methods. This integration provides a new perspective in understanding long-horizon videos. In fact, we propose a framework that allows for the integration of any neural perceptual models into our work, enhancing the capability to understand videos and enabling us to localize frames of interest with respect to queries. We show a comparison between different neural perceptual models in Tab. 2 and Fig. 5(a). We observe that NSVS-TL with various neural perception models performs differently depending on the complexity of the TL specification and datasets. The variation underscores the importance of these models’ inherent visual detection capabilities for scene identification.
@@
We evaluate TL-based reasoning against LLM-based reasoning over per-frame semantic annotations derived from neural perception models. The evaluation is conducted on four datasets: COCO-TLV, ImageNet-TLV, NuScenes-TLV, and Waymo-TLV. The results are collectively presented in Fig. 6. For single event scenarios such as “Always/Eventually Event A Occurs”, both our method and LLM-based reasoning perform reasonably well since these events do not require complex reasoning. In multi-event scenarios (“Event A and Event B”, “Event A until Event B," and “Event A and Event B until Event C"), our TL-based reasoning outperforms all LLM-based baselines.
@@
We evaluate multi-event sequences with temporally extended gaps, such as those in the specification of events A until B. For these scenarios, we design a scene identification task that extends the video length: event A starts at the beginning and continues until event B occurs at the end. This tests the reasoning capacity of the LLM benchmarks as the temporal distance between events increases. The scene identification performance of GPT-3.5 and GPT-3.5 Turbo Instruct degrade within videos that are $500$ seconds long as illustrated in Fig. 5(b). GPT-4, on the other hand, suffers from a sharp decline in performance for videos whose lengths range from $1000$ seconds and more. In contrast, NSVS-TL maintains consistent performance throughout. This consistency is observed even in videos spanning up to $40$ minutes or $2400$ seconds, indicating NSVS-TL’s potential reliability in handling even longer videos.
@@ 8 Conclusion
NSVS-TL, a neuro-symbolic method, enhances video understanding and scene identification in contexts demanding extended temporal reasoning. NSVS-TL, advances scene identification across various video durations by integrating semantic understanding with temporal reasoning. Unlike other methods focusing on short-duration video reasoning, we propose a novel approach for comprehending long-horizon, temporally extended videos.
@@
Acknowledgments
This material is based upon work supported in part by the Office of Naval Research (ONR) under Grant No. N00014-22-1-2254 and N00014-24-1-2097. Additionally, this work was supported by the Defense Advanced Research Projects Agency (DARPA) contract FA8750-23-C-1018 and DARPA ANSR: RTX CW2231110. Approved for Public Release, Distribution Unlimited.
@@ 9.1 Examples of Temporal Logic
Note that compared to LTL, CTL can quantify multiple branches while LTL is capable of one path. However, LTL can encode formulas like "Eventually-Always something happens," while CTL is not capable of.
@@
LTL over infinite traces was introduced in the field of computer science as a specification language designed for concurrent programs [23]. It facilitates the characterization of temporal reasoning within the domains of propositions and first-order logic. In our approach, frames of interest are considered as a finite set of frames extracted from a video. We can employ Linear temporal logic over finite traces ($\text{LTL}_{f}$) to handle finite sequences encompassing finite-length trajectories, frames from video, system procedures. Given a set of atomic propositions $P$, the syntax of $\textit{LTL}_{f}$ formulas is defined as follows:
@@ 9.2 LLM Benchmark Prompts
Prompt for LLMs You are a Language Assistant to enable searching a video frame for objects satisfying a condition. TASK DESCRIPTION: I will be giving you the list of objects detected in the video and a condition from the search. The list of objects will be specified in this format and include those frames where an object is detected: - Frame 1: [object1, object2, object3] - Frame 2: [object1, object2] A condition is of the form: - object1 Until object2, - object1 Until (object2 and object3), - Eventually object 1, - Always object 2, etc. You can use the list of objects detected in the video to help you. You don’t have access to the video. Your job is to specify the frame numbers where the rule is satisfied. There can be many such sequences. If you cannot find any sequence, you must return output [None]. To clarify, you will be looking for frames where the rule is satisfied. The output is a list of frame numbers, such as [1,20, 25, 50], specifying the frames where the rule holds. An example of the list of detections per frame is as follows: - 01: [person, car] - 02: [person, car] - 10: [person, car] - 30: [truck] - 40: [car] - 44: [person] - 50: [car, truck] Case 1: ‘person’ until ‘truck’ gives all frames of ‘person’ and ‘truck’ if ‘truck’ shows up.
Case 2: Eventually ‘person’ gives all frames of ‘person’ eventually showing up.
Case 3: ‘car’ and ‘truck’ should give all frames of ‘car’, and ‘truck’ showing up in the same frame.
Case 4: Always ‘car’ should give all frames of ‘car’ always showing up.
Case 5: (‘car’ and ‘truck’) until ‘person’ gives all frames of ‘car’ and ‘truck’ showing up in the same frame if ’person’ shows up; all frames of ’person’ included.